For the following experiments, compute the 95% confidence intervals and determine whether the data comes from the stated population:


(a) X= (3, 2.3, 4.5, 1.2, 5.6, 2.3, 4.5). =3.


(b) i = 4.2, ở = 1.2, n= 100, u= 3.


(c) ji = 0.2, ở = 5, n= 10, u= 3.


(d) å = 1.5, n=10, a=5.

1
Expert's answer
2021-12-01T16:38:56-0500

a)

"\mu=3.34, \sigma=1.45"


"t=\frac{x-\mu}{\sigma/\sqrt n}=\frac{x-3.34}{1.45/\sqrt 7}"


"df=n-1=6"

critical values:

"t_{crit}=\pm 2.447"


"-2.447<\frac{x-3.34}{1.45/\sqrt 7}<2.447"


"-1.34<x-3.34<1.34"

"2<x<4.68"

"x=3" comes from the stated population


b)

"t=\frac{i-u}{\sigma/\sqrt n}=\frac{i-3}{1.2/\sqrt {100}}"


"df=n-1=99"

critical values:

"t_{crit}=\pm 1.984"


"-1.984<\frac{i-3}{1.2/\sqrt {100}}<1.984"


"-0.02<i-3<0.02"

"2.98<i<3.02"

"i=4.2" does not come from the stated population


c)

"t=\frac{ji-u}{\sigma/\sqrt n}=\frac{ji-3}{5/\sqrt {10}}"


"df=n-1=9"

critical values:

"t_{crit}=\pm 2.262"


"-2.262<\frac{ji-3}{5/\sqrt {10}}<2.262"


"-3.58<ji-3<3.58"

"-0.58<ji<6.58"

"ji=0.2" comes from the stated population


d)

"t=\frac{x-a}{\sigma/\sqrt n}=\frac{x-5}{1.5/\sqrt {10}}"


"df=n-1=9"

critical values:

"t_{crit}=\pm 2.262"


"-2.262<\frac{x-5}{1.5/\sqrt {10}}<2.262"


"-1.07<x-5<1.07"

"3.93<x<6.07"

"a=5" ​comes from the stated population